# VeriLLM: A Lightweight Framework for Publicly Verifiable Decentralized Inference

### v3

@@ Abstract
Decentralized inference provides a scalable and resilient paradigm for serving large language models (LLMs), enabling distributed resource utilization and reducing reliance on centralized providers. However, in a permissionless environment without trusted nodes, ensuring the correctness of model outputs remains a core challenge. We introduce VeriLLM, a publicly verifiable protocol for decentralized LLM inference that achieves security under a one-honest-verifier assumption while maintaining practical efficiency. VeriLLM combines lightweight empirical rerunning with cryptographic commitments, allowing verifiers to validate results at approximately 1% of the underlying inference cost. To prevent verification bottlenecks, we design an isomorphic inference–verification architecture that multiplexes both inference and verification roles across the same GPU workers. This design (i) improves GPU utilization and overall throughput, (ii) enlarges the effective validator set, enhancing robustness and liveness, and (iii) enforces task indistinguishability to prevent node-specific optimizations or selective behavior. Through theoretical analysis and system-level evaluation, we show that VeriLLM achieves reliable public verifiability with minimal overhead, offering a practical foundation for trustworthy and scalable decentralized LLM inference.
1 Introduction
@@
The recent surge of Large Language Models (LLMs) is driven by their broad utility in diverse domains, including software development [wang2025enhancingcodellmsreinforcement] and creative content generation, resulting in widespread adoption across both industrial and academic settings [brown2020language, openai2023gpt4, touvron2023llama]. However, this progress exposes a fundamental weakness in current AI infrastructure: the computational and model resources required to support large-scale inference are heavily centralized within a few dominant technology providers [bender2021dangers, zaharia2023accelerating, sevilla2022compute]. Such concentration leads to systemic risks, including privacy violations, restricted access, single points of failure, and monopolistic control over core AI capabilities [bender2021dangers, whittlestone2021sociotechnical, weidinger2023sociotechnical, hazan2023foundation].
@@
To mitigate these risks, decentralized inference [yuan2022decentralized, morpheus2024, gonka2024, awsneuron2024, mei2024fedmoe, menon2024ravnest, yin2024knowledge, qu2025prompt] has emerged as a promising direction. By distributing computation across independent providers, decentralized systems aim to achieve openness, fault tolerance, and fairer access to AI resources [allen2022decentralized, kuppuswamy2022decentralizedai, raj2023federated, he2021trustworthy]. This paradigm enables verifiable collaboration among untrusted nodes and democratizes access to model execution, thereby improving transparency and resilience of large-scale AI deployments [villalobos2022future, bai2023open, kuppuswamy2022decentralizedai, zaharia2023accelerating].
@@
A central challenge in decentralized inference, however, is verifiable correctness. In a permissionless network, inference providers may be faulty, compromised, or economically motivated to skip computation. Without a mechanism to verify results, participants cannot ensure the integrity of model outputs [castro1999practical, cachin2011introduction, zhang2020delphi, xu2021verifynet, zamani2023decentralized]. This issue extends beyond adversarial behavior: providers often apply hidden optimizations—such as reduced numerical precision, pruned or distilled models, modified prompts, or compressed key–value caches—that degrade output fidelity [ho2018security, wang2022publiccheck, scaramuzza2025zkm, ong2025toploc, wang2024modelcompression, xu2023compressprompt]. As a result, independent verification becomes an operational necessity for any decentralized inference infrastructure [huang2024trustworthiness].
Existing approaches to verification face critical efficiency and deployability limitations. Cryptographic proof systems, including zero-knowledge proofs (ZKPs), can formally attest to correctness but remain computationally prohibitive: current frameworks introduce overheads several orders of magnitude larger than native inference, particularly for transformer-scale models [ezkl-github, chen2024zkml, sun2024zkllm, qu2025zkgpt]. Consensus-based schemes, in contrast, rely on redundant replication and majority voting or interactive verification games. These methods require strong synchrony or honest-majority assumptions and suffer from high latency and resource overhead [arun2025verde, mira-network, ambient-ai]. Furthermore, both proof-based and replication-based approaches align poorly with modern blockchain environments, where on-chain verification of large ML proofs is expensive and deterministic numerical reproducibility across heterogeneous hardware remains unsolved.
To address these limitations, we introduce VeriLLM, a publicly verifiable protocol for decentralized LLM inference designed for efficiency, security, and scalability. VeriLLM achieves verifiability under a one-honest-verifier assumption, enabling robust correctness guarantees without relying on honest-majority trust. The system combines lightweight empirical rerunning with cryptographic commitments, reducing the average verification overhead to about 1% of standard inference cost. In addition, a peer-auditing mechanism ensures collusion resistance and discourages lazy verification, enabling open, reliable, and low-cost validation of large-scale model execution.
To achieve security under the one-honest-verifier assumption while maintaining low verification cost, we design a hybrid architecture that integrates empirical rerunning with cryptographic verification. The proposed system consists of three core components that collectively ensure both efficiency and verifiability.
(1) Commitment of Intermediate States. During computation, all intermediate hidden states are committed to Merkle trees [merkle1989certified], with their root hashes recorded on-chain. These commitments yield tamper-evident logs, serving as binding evidence for subsequent verification and dispute resolution.
(2) Randomized Empirical Verification. A verifiable random function (VRF) [micali1999verifiable] determines unpredictable verification indices. At these positions, any verifier can re-execute the corresponding computations off-chain and compare the locally obtained results with the committed Merkle values. This sampling-based mechanism substantially reduces verification overhead while maintaining probabilistic soundness.
(3) Dispute-Resolution Protocol. In the event of inconsistencies, any honest verifier may initiate a dispute-resolution procedure by submitting Merkle proofs that demonstrate mismatches between recomputed and committed states. When empirical verification is insufficient for resolution, the protocol escalates to zero-knowledge proofs [ben1988knowledge, groth2016size] to provide conclusive adjudication.
Through this hybrid framework, our system ensures that even a single honest verifier can detect and prove computational misconduct with overwhelming probability, while keeping the expected verification cost low.
Building on this foundation, we further optimize the verification path by exploiting the inherent computational structure of large language model (LLM) inference. A standard inference pipeline consists of a Prefill phase followed by an autoregressive Decoder phase, with the latter dominating total runtime [agrawal2023]. We observe that when the complete output sequence is available to verifiers, the expensive Decoder phase need not be re-executed. Instead, verifiers can perform only the Prefill computation to validate correctness, as the subsequent token generation is a deterministic function of Prefill activations. This selective recomputation significantly reduces the verification overhead—approximately 1% of full inference cost—without compromising fidelity or verifiability.
Low verification cost alone is not sufficient in a permissionless network: validation must also be publicly auditable and resistant to non-participation. To achieve public verifiability and prevent lazy verification, VeriLLM augments the above optimization with a blockchain-assisted protocol that provides transparent coordination and accountable reporting, while preserving low on-chain cost. Intensive checks are executed off-chain by selected verifiers, and the blockchain performs lightweight consistency tests and resolves disputes. Sampling randomness is derived from on-chain verifiable random functions (VRFs) [micali1999verifiable], which select unpredictable hidden-state positions for audit. For each sample, verifiers submit Merkle openings [merkle1989certified] with minimal recomputation to attest that the required steps were executed. A peer-consistency mechanism then compares submissions and rewards alignment with the distribution expected from honest computation, which discourages collusion and maintains liveness without heavy replication.
Building on this foundation, we replace the conventional heterogeneous Inferencer–Verifier separation with an isomorphic inference–verification network in which all GPU workers multiplex both roles. The scheduler randomly assigns inference and verification jobs across the same hardware pool, and each worker executes both types without explicit role awareness. This task indistinguishability prevents behavior conditioned on job type, mitigates selective optimization, and improves hardware utilization by amortizing verification across the same computational substrate. As a result, the architecture simultaneously increases throughput and expands the effective validator set, strengthening robustness and decentralization.
Finally, VeriLLM undergoes comprehensive robustness evaluation against realistic threats such as quantized computation, early termination, and forged-output replay. The combination of Merkle commitments, VRF-based sampling, and on-chain dispute resolution ensures that any single honest verifier can detect and prove inconsistency with overwhelming probability. This establishes practical correctness guarantees without relying on replication or majority consensus.
1.1 Summary of Contributions
  • Single-Honest-Verifier Security: VeriLLM achieves correctness guarantees under a one-honest-verifier assumption through Merkle commitments, VRF-based sampling, and on-chain dispute resolution.
@@ 1.1 Summary of Contributions
  • Public and Collusion-Resistant Verification: A blockchain-based two-layer verification mechanism and peer-consistency incentives ensure that all verifiers perform computations honestly while keeping on-chain operations lightweight.
@@
  • Comprehensive Security Evaluation: We demonstrate robustness against practical attack vectors and show that VeriLLM provides scalable, low-cost, and publicly verifiable decentralized inference.
2 Design Goals and Threat Alignment
@@
This section formalizes the security and system design goals of VeriLLM and establishes their alignment with the assumed adversarial capabilities. Our design aims to provide verifiable, tamper-evident, and publicly auditable large language model (LLM) inference in a decentralized environment without relying on trusted majorities or centralized coordination [cachin2011introduction, allen2022decentralized, zhang2020delphi].
2.1 Adversarial Capabilities
@@
We consider a permissionless network composed of three types of participants: Inferencers, Verifiers, and a Scheduler. Each participant may deviate from the prescribed protocol to gain computational or economic advantage [castro1999practical, arun2025verde].
@@
  • C1: Malicious Inferencer. An Inferencer may attempt to reduce computation cost by (i) running a smaller or quantized model, (ii) terminating early, or (iii) reusing cached or fabricated hidden states while pretending to produce valid outputs [chen2024zkml, ezkl-github].
@@
  • C2: Lazy or Colluding Verifiers. A Verifier may skip recomputation and submit a verdict opportunistically, or collude with other verifiers or the Inferencer to falsely validate an incorrect computation [qu2025zkgpt, zamani2023decentralized].
@@
  • C3: Byzantine Scheduler. The Scheduler may selectively relay, alter, or censor messages; it may attempt to bias role assignment or falsify commitments to conceal misbehavior [ho2018security].
@@
  • C4: Observation and Timing Adversary. The adversary can observe network-level messages and timing patterns in an attempt to infer whether a node is serving inference or verification, and may adapt behavior based on such information [scion2015secure].
@@
  • C5: On-chain Adversary. Attackers may attempt to bias or replay randomness, forge signatures, or submit inconsistent commitments to gain advantage in the commit–reveal process [micali1999verifiable, bonneau2015sok].

### v4

@@ Abstract.
Decentralized inference provides a scalable and resilient paradigm for serving large language models (LLMs), enabling fragmented global resource utilization and reducing reliance on centralized providers. However, in a permissionless environment without trusted nodes, ensuring the correctness of model outputs remains a core challenge. We introduce VeriLLM, a publicly verifiable protocol for decentralized LLM inference that achieves security with incentive guarantees while maintaining practical efficiency. VeriLLM combines lightweight empirical rerunning with minimal on-chain checks to preclude free-riding, allowing verifiers to validate results at approximately 1% of the underlying inference cost by exploiting the structural separation between prefill and autoregressive decoding. To prevent verification bottlenecks, we design an isomorphic inference–verification architecture that multiplexes both inference and verification roles across the same GPU workers. This design (i) improves GPU utilization and overall throughput, (ii) enlarges the effective validator set, enhancing robustness and liveness, and (iii) enforces task indistinguishability to prevent node-specific optimizations or selective behavior. Through theoretical analysis and system-level evaluation, we show that VeriLLM achieves reliable public verifiability with minimal overhead, offering a practical foundation for trustworthy and scalable decentralized LLM inference.
Decentralized Inference, Verification, Blockchain, Game Theory
copyright: noneconference: arxiv; 2026; Onlineccs: Security and privacy Distributed systems securityccs: Security and privacy Security protocolsccs: Computer systems organization Peer-to-peer architectures
1. Introduction
@@
The recent surge of Large Language Models (LLMs) is driven by their broad utility in diverse domains, including software development (Wang et al., 2025) and creative content generation, resulting in widespread adoption across both industrial and academic settings (Brown et al., 2020; OpenAI, 2023; Touvron et al., 2023). However, this progress exposes a fundamental weakness in current AI infrastructure: the computational and model resources required to support large-scale inference are heavily centralized within a few dominant technology providers (Bender et al., 2021; Zaharia et al., 2023; Sevilla et al., 2022). Such concentration leads to systemic risks, including privacy violations, restricted access, single points of failure, and monopolistic control over core AI capabilities (Bender et al., 2021; Whittlestone et al., 2021; Weidinger et al., 2023; Hazan et al., 2023).
@@
To mitigate these risks, decentralized inference (Yuan et al., 2022; Network, 2024a; Foundation, 2024; Amazon Web Services, 2024; Mei et al., 2024; Menon et al., 2024; Yin et al., 2024; Qu et al., 2025c) has emerged as a promising direction. By distributing computation across independent providers, decentralized systems aim to achieve openness, fault tolerance, and fairer access to AI resources (Allen et al., 2022; Kuppuswamy and Shetty, 2022; Raj et al., 2023; He et al., 2021). This paradigm enables verifiable collaboration among untrusted nodes and democratizes access to model execution, thereby improving transparency and resilience of large-scale AI deployments (Villalobos et al., 2022; Bai et al., 2023; Kuppuswamy and Shetty, 2022; Zaharia et al., 2023).
@@
A central challenge in decentralized inference, however, is verifiable correctness. In permissionless environments, providers are economically motivated to skip computation (”lazy workers”) or apply hidden optimizations—such as quantization or pruning—that degrade output fidelity (Ho et al., 2018; Wang et al., 2022; Scaramuzza et al., 2025; Ong et al., 2025; Wang et al., 2024; Xu et al., 2023) Without a robust verification mechanism, participants cannot ensure the integrity of model outputs (Castro and Liskov, 1999; Cachin et al., 2011; Zhang et al., 2020; Xu et al., 2021; Zamani and Shokri, 2023). As a result, independent verification becomes an operational necessity for any decentralized inference infrastructure (Huang et al., 2024).
At a fundamental level, verifiable LLM inference in permissionless environments faces inherent tensions among computational efficiency, numerical fidelity, and economic incentives. Exact verification is infeasible due to floating-point nondeterminism and hardware heterogeneity, while replication-based approaches incur prohibitive cost and collapse under rational laziness (Verifier’s Dilemma). Cryptographic proofs (e.g., ZK-SNARKs) formally attest to correctness but incur overheads several orders of magnitude higher than native inference, rendering them impractical for transformer-scale models (ZKonduit, 2025; Chen et al., 2024; Sun et al., 2024; Qu et al., 2025b). Consensus-based approaches rely on replication and voting but suffer from critical scalability flaws: they incur multiplicative computational costs, struggle with floating-point non-determinism across heterogeneous devices, and require an impractically large set of honest nodes to maintain security guarantees. Furthermore, both approaches align poorly with modern blockchain environments, where on-chain verification of large ML proofs is economically infeasible and numerical reproducibility across heterogeneous hardware remains unsolved.
To address these limitations, we introduce VeriLLM, a high-assurance protocol for decentralized inference that achieves publicly verifiability with negligible overhead. By strategically leveraging the structural asymmetry of LLM execution, VeriLLM reduces the verification burden to approximately 1% of the inference cost. Crucially, unlike consensus protocols that demand a synchronous honest majority within every local committee, VeriLLM guarantees security relying only on a relaxed global honest majority assumption. This trust model is realized through a single-round re-verification mechanism. To further harden this foundation, we design a task-indistinguishable isomorphic network architecture where nodes dynamically multiplex roles. This design maximizes the effective validator set and significantly enhances security against selective attacks and resource efficiency. Furthermore, VeriLLM employs a sampling-based on-chain comparison mechanism that simultaneously resolves the Verifier’s Dilemma (lazy verification) and enables lightweight on-chain settlement, while remaining robust to floating-point non-determinism.
We minimize verification costs by leveraging a structural asymmetry in Transformer execution. Since the verifier already possesses the complete output sequence claimed by the worker, it can bypass the expensive, sequential autoregressive decoding process. Instead, the verifier concatenates the input prompt with the generated output tokens and processes the entire sequence as a single, fully parallelizable Prefill operation. This shifts the workload from a memory-bandwidth-bound serial process to a compute-bound parallel operation that efficiently saturates GPU Tensor Cores. Consequently, verification overhead drops to approximately 1% of the inference cost—an advantage that becomes even more pronounced in long-context tasks, where the marginal cost of parallel verification is negligible.
To decisively resolve the Verifier’s Dilemma, we enforce a sampling-based audit mechanism. Verifiers are required to submit not only their final judgment but also specific sampled hidden states. These lightweight state fragments undergo on-chain adjudication via the smart contract, where any deviation triggers immediate penalization. To ensure comprehensive security, we design a distinct comparison protocol tailored to different adversarial scenarios. To address floating-point non-determinism, we introduce a noise-tolerant comparison algorithm specifically optimized for floating-point arithmetic. By experimentally calibrating acceptance thresholds based on real-world hardware, this algorithm can effectively filter out numerical divergence, preventing false positives while maintaining strict sensitivity to malicious deviations.
While VeriLLM operates under a global honest-majority assumption, a critical challenge is the risk of ’local majority failure’, where a randomized committee happens to be captured by a coordinated adversary. To mitigate this, we incorporate a re-verification scheme that allows disputes to be escalated to a newly sampled set of verifiers. A central finding of our work is that a single round of re-verification is mathematically sufficient to secure the protocol. Our game-theoretic analysis confirms that the mere option of this one-time escalation effectively anchors local decisions to the global honest majority; this renders honest voting the unique Nash Equilibrium, guaranteeing robustness without incurring the latency of recursive arbitration.
To guarantee protocol integrity without centralized trust, VeriLLM incorporates a suite of cryptographic primitives. First, we employ Verifiable Random Functions (VRFs) to govern both node selection and hidden-state sampling. This ensures that randomness is publicly verifiable and unpredictable, thereby preventing the scheduler from biasing the audit process. Second, to secure on-chain data submission, we implement a multi-stage commit-reveal scheme. This mechanism forces nodes to bind their results before revealing them, effectively neutralizing ’copycat’ attacks (front-running) and preventing ex-post facto modification. Finally, we leverage Merkle Proofs to anchor off-chain data, enabling succinct and gas-efficient on-chain verification of data existence and integrity.
Contributions. We make the following contributions:
@@
  • Global Honest Majority Requirement: Our design only needs an honest majority among global verification nodes instead of committee nodes, enhancing robustness in verification processes.
  • Public and Collusion-and-Lazy-Resistant Verification: A blockchain-based two-layer verification mechanism ensure that all verifiers perform computations honestly while keeping on-chain operations lightweight. This design also bypasses the Verifier’s Dilemma via a peer-consistency framework (Zhao et al., 2024) that disincentivizes lazy verification.
@@
  • Noise-Resistant Design: Our protocol ensures correctness and soundness even in the presence of machine and floating-point errors.
@@
  • Comprehensive Security Evaluation: We theoretically demonstrate robustness against practical attack vectors and show that VeriLLM provides scalable, low-cost, and publicly verifiable decentralized inference. We implement VeriLLM, and our empirical results substantiate our theoretical claims.
2. Related Work
2.1. Activation Sampling & Compact Commitment
Lightweight methods verify inference integrity by selectively committing to critical activation states. TOPLOC/INTELLECT (Ong et al., 2025) commits to the top-128 final-layer activations every 32 tokens, compressing data 1000× via polynomial congruence (258 bytes per commitment). However, its verification process requires a trusted third party to re-run the computation and provide a ground truth, which reintroduces a central trust assumption. TensorBlock (TensorBlock, 2025) extends this to KV-cache sampling using deterministic token/layer selection, enabling validators to recompute partial trajectories from checkpoints. VeriSplit (Zhang et al., 2024) supports private outsourcing through linear operator masking: clients add noise to inputs, workers return encrypted activations, and clients locally denoise results using precomputed terms, with Merkle trees enabling partial verification. However, their probabilistic security cannot guarantee the detection of sophisticated model tampering, and input privacy relies on noise injection techniques that are vulnerable to advanced reconstruction attacks.
2.2. Zero-Knowledge Machine Learning (ZKML)
Sumcheck-based Protocols: zkCNN (Liu et al., 2021) accelerates convolution verification by transforming operations to the frequency domain. VerfCNN (Qu et al., 2025a) achieves theoretical optimal prover performance in convolution verification. zkLLM (Sun et al., 2024) reduces communication by 90% for non-arithmetic operations (e.g., Softmax) via the tlookup protocol and exploits translation invariance. QAP-based Protocols: vCNN (Lee et al., 2024) represents convolutions as Quadratic Polynomial Programs (QPP), reducing multiplicative gates exponentially. pvCNN (Weng et al., 2023) further optimizes tensor operations via Quadratic Matrix Programs (QMP) and batch verification. ZKML (Chen et al., 2024) offers modular circuit design with 43 pre-built layer circuits and automated optimization. zkGPT (Qu et al., 2025b) leverages multi-thread parallelization for LLM inference-proof acceleration. Lookup arguments (Campanelli et al., 2024) enable matrix-level privacy with 30% faster proofs. VOLE-based Protocols: Mystique (Weng et al., 2021) achieves sublinear communication for matrix multiplication using sVOLE-based tensor compression. Recent frameworks (Hao et al., 2024) optimize nonlinear functions via table lookups and digital decomposition, reducing constraints by over 99%.
System & Hardware Optimization: zkLoRA (Roy et al., 2025) supports verifiable fine-tuning via incremental proofs for LoRA modules. EZKL (ZKonduit, 2025) automates ZKP circuit generation from ONNX models. DeepProve (Lagrange Labs, 2025) accelerates GPU-based ZKML (54–158× faster proofs via parallelized SNARKs). Sertn (Inference Labs, 2025) employs optimistic ”Staked Deferred Proofs” with economic slashing for invalid results. However, these methods impose prohibitive prover costs (hours per inference), require trusted setups, and suffer from rigidity in supporting novel architectures due to circuit specialization.
2.3. Trusted Execution Environments (TEEs)
Atoma Network (Atoma Network, 2025) combines TEE attestation with sampling-based consensus for deterministic execution. evML (Exo Explore, 2025) uses TEE-based remote attestation (e.g., Google/Apple frameworks) with probabilistic auditing. nesa.ai (NESA, 2025) orchestrates heterogeneous TEEs for secure multi-party inference. Phala Network (Network, 2024b) extends this paradigm by providing a ”Confidential AI” cloud platform that enables standard Dockerized AI applications to run within GPU-equipped TEEs. However, they inherit vulnerabilities to side-channel exploits (e.g., cache-timing attacks), centralize trust in proprietary hardware (Intel SGX/ARM TrustZone), and exhibit limited auditability across heterogeneous environments.
2.4. Consensus & Optimistic Approaches
Verde (Arun et al., 2025) employs optimistic dispute resolution, recomputing only divergent operators (e.g., single attention heads) to minimize overhead. Mira (Mira, 2025) leverages multi-model consensus against hallucinations using domain-specific thresholds. Ambient (Ambient.ai, 2025) introduces Proof-of-Logits (PoL), hashing intermediate logits for spot-check verification at $0.1\%$ overhead. However, they introduce delayed finality during dispute windows and implicitly assume honest majority participation.
3. Design Goals and Threat Alignment
@@
This section formalizes the security and system design goals of VeriLLM and establishes their alignment with the assumed adversarial capabilities. Our design aims to provide verifiable, tamper-evident, and publicly auditable large language model (LLM) inference in a decentralized environment without relying on trusted majorities or centralized coordination (Cachin et al., 2011; Allen et al., 2022; Zhang et al., 2020).
3.1. Adversarial Capabilities
@@
We consider a permissionless network composed of three types of participants: Inferencers, Verifiers, and a Scheduler. Each participant may deviate from the prescribed protocol to gain computational or economic advantage (Castro and Liskov, 1999; Arun et al., 2025).
@@
  • C1: Malicious Inferencer. An Inferencer may attempt to reduce computation cost by (i) running a smaller or quantized model, (ii) terminating early, or (iii) reusing cached or fabricated hidden states while pretending to produce valid outputs (Chen et al., 2024; ZKonduit, 2025).
@@
  • C2: Lazy or Colluding Verifiers. A Verifier may skip recomputation and submit a verdict opportunistically, or collude with other verifiers or the Inferencer to falsely validate an incorrect computation (Qu et al., 2025b; Zamani and Shokri, 2023).
@@
  • C3: Byzantine Scheduler. The Scheduler may selectively relay, alter, or censor messages; it may attempt to bias role assignment or falsify commitments to conceal misbehavior (Ho et al., 2018).
@@
  • C4: Observation and Timing Adversary. The adversary can observe network-level messages and timing patterns in an attempt to infer whether a node is serving inference or verification, and may adapt behavior based on such information (Perrig et al., 2015).
@@
  • C5: On-chain Adversary. Attackers may attempt to bias or replay randomness, forge signatures, or submit inconsistent commitments to gain advantage in commit–reveal process (Micali et al., 1999; Bonneau et al., 2015).